# When recalling in-context, Transformers are not SSMs

### v1

@@ 3 Closer Look into AR performance
Building on previous research, we aim to provide an in-depth analysis of the differences and similarities between attention and recurrent models through the lens of AR. Prior studies (Arora et al., 2023) have shown that transformers are inherently well-suited for solving the MQAR task, achieving perfect accuracy regardless of model dimension, sequence length or number of key-value pairs to infer. In contrast, it was argued (both theoretically and empirically) that new recurrent models (Peng et al., 2023; Nguyen et al., 2024; Gu and Dao, 2024) can only solve MQAR if the hidden dimension is roughly equal to the sequence length (see analysis by Jelassi et al. (2024) in a related setting). However, a key aspect that has been overlooked in some prior works is the crucial role of optimization in recurrent models —particularly, the use of an effective grid search for the choice of learning rate.
Hypothesis from previous works.
Recurrent models update their hidden state (which serves as a compressed representation of past information) at each time step, using the current input. Since the model only has access to its hidden state and the current input, its ability to recall previous information depends on how effectively it compresses past data into this state. With a simplified analysis assuming uniform distribution over strings, Jelassi et al. (2024) showed that to successfully copy input strings, the hidden size needed grows linearly with the sequence length. In contrast, transformers  (Vaswani et al., 2017) dynamically access all previously seen inputs through the softmax attention mechanism, allowing for the explicit computation of interactions between tokens. This makes the task of recalling already seen tokens essentially a lookup table problem when two layers work simultaneously, as described in Jelassi et al. (2024); Olsson et al. (2022).
Figure 2: Performance of 2-layers models. Results for H3 (Fu et al., 2023), RWKV (Peng et al., 2023) and Based (Arora et al., 2024) included in App. A.3. We report the official results [^3] Mamba was not included in the official work but some experiments are documented in the blog post ( green stars) and the replication running the original code of (Arora et al., 2023) ( dotted blu line). While for replication, we used the learning rates grid by Arora et al. (2023), we note here that, due to high sensitivity to the learning rate (Fig, 1), tuning drastically affects performance. In solid orange, we provide results with a finer grid (cf. Fig.1). Careful tuning of the learning rate gives a general improvement in the performance of recurrent models. This becomes especially crucial in Mamba, where the task becomes solvable at high sequence lengths >>>> hidden size. The results show the mean and relative max-min errors for 3 seeds. Attention always solves the task (all curves overlap).
@@ Results.
Compared to previous work, in our experiments, we devoted more attention to tuning the learning rates, drastically improving the reported performance for recurrent models (see Fig. 3&1). As shown in Figure 3 and extensively in Appendix A.3, a finer grid not only enhances average performance across all settings but also proves particularly crucial for the Mamba model. With a more suitable learning rate, Mamba (Gu and Dao, 2024), which was previously shown to struggle with long sequence lengths, becomes capable of solving MQAR at relatively small hidden model sizes. All experimental details for this and the next experiments are in Appendix A.2. This highlights a key takeaway for MQAR: the choice of learning rate (and optimization strategy in general) can be decisive in assessing whether a recurrent model can solve the task at all. In the case of Mamba, optimization choices become a discriminative factor, emphasizing the necessity of careful hyperparameter tuning in recurrent models, and further research for improving their high sensitivity.
To further emphasize the critical role of learning rate selection in training recurrent models, we compare the performance of Attention, Hyena and Mamba using the same grid search. Figure 1 illustrates that attention-based models maintain strong performance across a relatively wide range of learning rates. In contrast, Hyena and Mamba exhibit a different behavior: performance remains near zero for most learning rates but suddenly reaches near-optimal levels at specific values which may not be included in the grid by Arora et al. (2023). These findings highlight a key distinction between attention-based and recurrent models: a sparse learning rate grid search can disproportionately impact their training outcomes. This discrepancy can lead to misleading conclusions about the capabilities of these models, emphasizing the need for careful tuning.
@@ 4 Effects of width/depth scaling into AR performance
Figure 3: Performance of 1-layer attention-based (Attention, Based) and recurrent-based (H3, Hyena, Mamba, RWKV) models on AR. We show how for recurrent models, scaling the width boosts performances. On the contrary, attention models cannot solve the task anymore as in the 2-layer setting, and performances are unaffected by the scaling in width. The results show the mean and relative max-min errors after 3 runs with different seeds.
While our findings in Sec. 3 show that some recurrent models can exhibit improved performance on MQAR with proper learning rate tuning, we confirm that a sizeable gap with attention can still be observed for some recurrent models at low widths (e.g. Hyena vs. Attention). The experiments of Sec. 3 focused on comparisons of 2-layer architectures, at different sequence lengths and model widths. This choice stems from prior research (Olsson et al., 2022), where transformers have shown peculiar in-context learning capabilities related to the formation of induction head circuits in 2-layer models. With the intention of going beyond the setup that is known to show strengths for softmax attention, our objective in this section is to explore the effects of scaling in different configurations.
To achieve this goal, we conducted experiments analogous to Section 3 using single-layer architectures [^4] By single layer in attention and recurrent models we mean a sequence mixer followed by an MLP.. By doing so, we aim to decouple the effects of inter-communication between layers and to isolate the impact of each model’s fundamental structure (attention versus recurrence) on MQAR. Beyond this, our motivation also comes from the notable connections that have been drawn between attention and recurrent models (Dao and Gu, 2024; Ali et al., 2024; Sieber et al., 2024) and on the capabilities of transformers  (Sanford et al., 2024) – all of which concern 1-layer models. Our results, presented in Figure 3, reveal two key insights:
@@
    Attention models exhibit a surprisingly different behavior: when constrained to a single layer, they fail to solve the task and increasing the hidden dimension does not affect their performance. This is in stark contrast to their strong results in 2-layer architectures, where even the smallest model was sufficient to solve the task in the hardest setting. Interestingly, in this setting transformers are capable on average of recalling one key-value pair in every setting, suggesting a memory size issue when only one layer is present as also suggested in previous work  (Sanford et al., 2024).
Figure 4: Scaling models in width and depth (Seq len: 256, KV pairs: 64). Symbols with the same shape and color represent models of increasing size in the following order: 64, 128, 256, 512, 1024, and 2048. We show how rather than the number of parameters, is the way these models are scaled that impacts performance. Specifically, recurrent models benefits from scaling in width, while attention benefits from scaling in depth.
@@
Figure 5: Training (lower opacity) and Validation dynamics of 1-layer models. We reported within brackets the smallest width that solves the task, if possible; or otherwise the biggest width we tried (for attention). Differently from Mamba, Hyena requires the model dimension to exceed the sequence length. Both exhibit smooth learning dynamics, leading to perfect performance. Attention shows a loss bump, but without accuracy gains, suggesting an attempt to form induction heads that the single-layer transformer fails to leverage effectively.
Sec. 4 sparked our curiosity, leading us to explore the single-layer architecture setup further – to understand why attention hits a performance ceiling while recurrent models can solve the task. This analysis is especially intriguing given the strong connections that have been proposed between attentions and Mamba in Ali et al. (2024); Dao and Gu (2024).
In this section, we analyze the training dynamics of well-tuned Hyena, Attention and Mamba models. As illustrated in Fig. 5 we identify two main patterns. First, Hyena (and similarly other non-selective recurrent models like H3 and RWKV) exhibits consistently smooth learning dynamics, with a gradual and steady improvement that eventually lead to convergence at the solution. Specifically, loss reductions align closely with increases in accuracy. Differently, attention accuracy remains largely unchanged throughout training. A similar trend appears in the test loss, which remains relatively stable until a sudden bump occurs, after which the test loss settles again. This bump resembles the formation of an induction head circuit (Olsson et al., 2022), and to the best of our knowledge has previously only been observed during the training of multi-layer transformer architectures. However, as opposed to what can observe in 2-layer models, this phase transition in the loss does not correspond to an accuracy improvement for attention. Based on previous work  (Olsson et al., 2022), we hypothesize that during this phase transition, the attention mechanism attempts to form induction heads. However, in the single-layer setting, the model lacks the expressivity needed to effectively leverage this mechanism for task resolution. Interestingly, the dynamics of Mamba is mixed:
@@
    Like single-layer attention models, we report a significant loss bump, reinforcing the connection between Mamba and attention mechanisms, as suggested in  Ali et al. (2024); Dao and Gu (2024).
@@
Our results highlight a crucial distinction: while attention and recurrent models share some common ground, yet distinct inductive biases. Moreover, their performance is in strong interaction with the optimization algorithm at hand (in our case, Adam (Kingma, 2014)), as we also saw in Figure 1. Understanding these nuances is key to optimally leveraging both architectures, perhaps also towards hybrid models (Waleffe et al., 2024; Dao and Gu, 2024).
6 Are SSMs and Transformers really similar?
Our results so far highlight key differences between Transformers and SSMs, particularly Mamba, in the context of AR. Notably, while Mamba demonstrates greater expressivity—successfully solving the task even in a single-layer setting—it presents optimization challenges in terms of learning rate stability. In contrast, Transformers exhibit remarkable stability across a wide range of suitable learning rates during training in the 2-layer setting. We highlight that each of these layers includes an MLP block processing channels.
To address this discrepancy, we conduct a series of ablation studies aimed at:
@@ 6 Are SSMs and Transformers really similar?
    Modifying the architectures of both models to better align them (architectures are in appendix A.1) and identify the source of Mamba’s superior performance, summarized in Table 1 and
@@
Table 1: Performance of ablated 1-layer architectures. After making different modifications to the Mamba and Transformer, we report if the model is capable of solving the MQAR task. Here solving the task means achieving accuracy $\geq$ 95 percent with any combination of sequence length 64, 128,2 56, 512 and model dimension 64, 128, 256, 512, 1024, 2048
Model Solves MQAR
Attention ×
Attention + Conv on QKV
Attention + Conv on K
Attention + Conv on V
Attention + Conv on Q ×
Mamba
@@
S6 + MLP (Mamba as a Transformer)
Convolutions.
Inspired by  (Li et al., 2024), we begin by adapting the attention mechanism to resemble Mamba. We incorporate a 1D convolution before the Query,Key and Value matrix projections to brings in locality, enabling the model to solve MQAR with just one layer. Interestingly, we observe that applying the convolution to either the Key or Value matrix alone is sufficient to achieve the same performance gains. These observations suggest that the 1D convolution may be a central factor behind Mamba’s effectiveness. However, we find that even after removing the convolution, the model retains the ability to solve the task.
Backbone ablation.
We further modify the Mamba architecture to make it closer to the Transformer architecture by: (1) removing the gating mechanism, and (2) replacing the standard Mamba block with the individual sequence mixer S6, followed by an MLP—mirroring the Transformer’s architecture. Despite these alterations, Mamba continues to perform well when properly tuned, suggesting a notable degree of robustness in the design of its fundamental block.
@@
To better understand what contributes to training stability, we also evaluate architectural variants designed for improving the Mamba architecture and solve the MQAR task. In particular, we test Mamba2  (Dao and Gu, 2024) and DeltaNet  (Yang et al., 2024b) as shown in Figure 6. While performance of Mamba2 is slightly more stable, Transformer-level robustness is only achieved by DeltaNet. A closer look at the DeltaNet update rule reveals that its mixing is based on Householder matrices. As such, the off-diagonal terms such as $C_{N}\prod_{k=1}^{N}A_{k}B_{0}$ do not necessarily incur in vanishing gradients. Instead, in both Mamba and Mamba2, $A_{k}$ includes a decay rate that induces vanishing gradients and fast decay of off-diagonal terms, as recently pointed out by Trockman et al. (2024). We hypothesize this is the main distinction unlocking stable optimization in DeltaNet.
Figure 6: We show the performance of Mamba, Mamba2 and DeltaNet in the 1-layer setting using the same learning rate grid search. Here we show how having a bigger hidden state marginally helps stability, as in Mamba2 and especially in DeltaNet. Our results are presented for model dimensions up to 256, which was the maximum size supported by the DeltaNet implementation. The results show the mean and relative max-min errors after 3 runs with different seeds.
@@ 7 Discussion and Conclusions
In this work, we used MQAR as a benchmark to compare attention and recurrent models at a small scale. Our findings shed additional light on how the underlying mechanisms of these models influence their performance. Specifically, we showed that recurrent models are highly sensitive to optimization, with their performance significantly affected by the choice of learning rate. This underscores the need for further research to improve their stability. Additionally, we observed contrasting scaling behaviors: recurrent models benefit from the increased width and hidden state size, whereas transformers struggle with MQAR in a single-layer configuration. Interestingly, despite their poor performance, single-layer transformers exhibit training dynamics resembling the induction head phenomenon, previously reported only in multi-layer settings. Instead, Mamba displays similar behavior but successfully solves the task. Finally, through the ablations study, we showed how the performance of Mamba is robust to specific architectural components such as gating and convolution, and how other similar architectures can enhance performance and stability. Our findings suggest overlaps between the optimization landscapes of Mamba and Attention, yet with crucial differences related to expressivity, to study further. Looking ahead, we think that exploring other synthetic reasoning tasks and architectural changes could provide further insights into the mechanisms behind these models. Evaluating these areas is an important direction to refine our understanding of modern sequence models.

### v2

@@ 3 Closer Look into AR performance
Building on previous research, we provide an in-depth analysis of the differences between attention-based and recurrent models through the lens of AR. Prior studies (Arora et al., 2023) have shown that Transformers are inherently well-suited for solving the MQAR task, achieving perfect accuracy across all settings. In contrast, it was argued (both theoretically and empirically) that new recurrent models can only solve MQAR if the hidden dimension is roughly equal to the sequence length. However, a key aspect that has been overlooked is the crucial role of optimization in recurrent models – particularly, the use of an effective grid search for the choice of learning rate.
The memory bottleneck hypothesis.
Recurrent models update their hidden state (which serves as a compressed representation of past information) at each time step, using the current input. Since the model only has access to its hidden state and the current input, its ability to recall previous information depends on how effectively it compresses past data into this state. For instance, with a simplified analysis assuming uniform distribution over strings, Jelassi et al. (2024) showed that to successfully copy input strings, the hidden size needed grows linearly with the sequence length. In contrast, Transformers dynamically access all previously seen inputs through the softmax attention mechanism, allowing for the explicit computation of interactions between tokens. This makes the task of recalling already seen tokens essentially a lookup table problem when two layers work simultaneously, as described by Jelassi et al. (2024); Olsson et al. (2022).
Figure 2: Performance of 2-layers models in MQAR. We report the official results [^4] Mamba was not included in the official work but some experiments are documented in the blog post. ( green stars) and the replication running the original code of (Arora et al., 2023) ( dotted blue line). While for replication, we used the learning rate grid by Arora et al. (2023), we note here that, due to high sensitivity to the learning rate (Fig, 1), tuning drastically affects performance. In solid orange, we provide results with a finer grid (cf. Fig.1). Careful tuning of the learning rate gives a general improvement in the performance of recurrent models. This becomes especially crucial in Mamba, where the task becomes solvable at high sequence lengths >>>> hidden size. The results show the mean and relative max-min errors for 3 seeds. Attention always solves the task (all curves overlap).
@@ Results.
Compared to previous work, in our experiments, we carefully tuned the learning rates, drastically improving the reported performance for recurrent models (see Fig. 4&1). As shown in Figure 4 (full tables in Appendix A.3 with more models), a finer grid not only enhances average performance across all settings but is also particularly crucial for the Mamba (Gu & Dao, 2024) model. With a more suitable learning rate, Mamba, which was previously shown to struggle with long sequence lengths, becomes capable of solving MQAR at relatively small hidden model sizes. All experimental details for this and subsequent experiments are in Appendix A.2. This highlights a key takeaway for MQAR: the choice of learning rate (and optimization strategy in general) can be decisive in assessing whether a recurrent model can solve the task at all. In the case of Mamba, optimization choices become a discriminative factor, emphasizing the necessity of careful hyperparameter tuning in recurrent models, and further research for improving their high sensitivity.
To further emphasize the critical role of learning rate selection in training recurrent models, we compare performance with respect of our grid search. Figure 1 illustrates that Attention-based models maintain strong performance across a relatively wide range of learning rates. In contrast, Hyena and Mamba exhibit a different behavior: performance remains near zero for most learning rates but suddenly reaches near-optimal levels at specific values that may not be included in the grid by Arora et al. (2023). These findings highlight a key distinction between Attention-based and recurrent models: a sparse learning rate grid search can disproportionately impact their training outcomes. This discrepancy can lead to misleading conclusions about the capabilities of these models, emphasizing the need for careful tuning.
@@ 4 Effects of width/depth scaling into AR performance
Figure 3: Performance of 1-layer models on MQAR. We show how for recurrent models, scaling the width boosts performance. On the contrary, Attention models can no longer solve the task anymore as they do in the 2-layer setting, and performances are unaffected by the scaling in width. The results show the mean and relative max-min errors after 3 runs with different seeds.
Figure 4: Scaling behavior (Seq len: 256, KV pairs: 64). Symbols with the same shape and color represent models of increasing dimension from 64 to 2048. We show that the scaling strategy, rather than the total number of parameters, is what primarily impacts performance. Specifically, recurrent models benefit from scaling in width, while attention-based models benefit from scaling in depth.
While our findings in Sec. 3 show that some recurrent models can exhibit improved performance on MQAR with proper learning rate tuning, we confirm that a sizable gap with Transformers can still be observed at low widths (e.g. Hyena). The experiments of Sec. 3 focused on comparisons of 2-layer architectures, at different sequence lengths and model widths. This choice stems from prior research (Olsson et al., 2022), where Transformers have shown peculiar in-context learning capabilities related to the formation of induction head circuits in 2-layer models. Indeed increasing the number of layers to more than 2 does not provide any further improvement in MQAR performance. With the intention of going beyond the setup that is known to show strengths for softmax attention, our objective in this section is to explore the effects of scaling in different configurations.
To achieve this, we conducted experiments analogous to Sec. 3 using single-layer models [^5] In this context, a single layer refers to a sequence mixer followed by an MLP.. By doing so, we aim to decouple the effects of inter-communication between layers and to isolate the impact of each model’s fundamental structure. Beyond this, our motivation also comes from the notable connections that have been drawn between Attention and recurrent models (Dao & Gu, 2024; Ali et al., 2024; Sieber et al., 2024) and on the capabilities of Transformers (Sanford et al., 2024) – all of which concern 1-layer models. Our results, presented in Figure 3 (full table in Appendix A.4), reveal two key insights:
@@
    Attention exhibit a surprisingly different behavior: when constrained to a single layer, they fail to solve the task and increasing the hidden dimension does not affect their performance. This is in stark contrast to their strong results in 2-layer architectures, where even the smallest model was sufficient to solve the task in the hardest setting. Interestingly, in this setting Transformers are capable on average of recalling one key-value pair in every setting, suggesting a memory size issue when only one layer is present as also suggested in previous work  (Sanford et al., 2024).
@@
5 Copy task
To validate that optimization instabilities and scaling behaviors observed on MQAR are not dataset-specific, we conducted a parallel investigation on the copying task studied by (Jelassi et al., 2024).
Figure 5: Performance of a Transformer with RoPE and Mamba on the copy task following  (Jelassi et al., 2024) implementation. This task also highlights the narrow window of suitable learning rates that allows Mamba to solve the task.
Learning Rate Instability.
Just as with MQAR, optimization stability offers a new perspective on the conclusions of (Jelassi et al., 2024). As shown in Figure5, the Transformer solves the copying task robustly across a wide range of learning rates. In contrast, Mamba’s success is again confined to a narrow window.
Parameter Scaling.
We find that attempts to provide fair comparisons by matching parameter counts through increased depth in SSMs are misguided. As shown in Table 1, a deeper but narrower Mamba fails to copy, whereas a shallower but wider Mamba with the same parameter count succeeds. This reinforces our claim that architectures must be scaled along their preferred axes – width for SSMs and depth for Transformers – to unlock their potential.
Table 1: Performance on the copy task. When comparing Transformers and SSMs with the same number of parameters, it is crucial to scale the latter in width rather than depth.
Architecture # Layers Width # Parameters (M) Accuracy (%)
Attention (RoPE) $12$ $1024$ $150$ $100\%$
Mamba $12$ $1024$ $80$ $0\%$
Mamba $24$ $1024$ $150$ $16\%$
Mamba $12$ $1408$ $150$ $100\%$
@@ 5 Copy task
Figure 6: Training (lower opacity) and validation dynamics of 1-layer models in MQAR. We report within brackets the smallest width that solves the task, if possible; or otherwise the largest width we tried (for Attention). Differently from Mamba, Hyena requires the model dimension to exceed the sequence length. Both exhibit smooth learning dynamics, leading to perfect performance. Attention shows a loss bump, but without accuracy gains, suggesting an attempt to form induction heads that a single-layer Transformer fails to leverage effectively.
Sec. 4 and  5 sparked our curiosity, leading us to explore the single-layer architecture setup further in MQAR – to understand why Attention hits a performance ceiling while recurrent models can solve the task. This analysis is especially intriguing given the strong connections that have been proposed between attention and Mamba in Ali et al. (2024); Dao & Gu (2024); Sieber et al. (2024).
In this section, we analyze the training dynamics of well-tuned Hyena, Attention and Mamba models. As illustrated in Fig. 6 we identify two main patterns. First, Hyena exhibits consistently smooth learning dynamics, with a gradual and steady improvement that eventually leads to convergence at the solution. Specifically, loss reductions align closely with increases in accuracy. Differently, Attention accuracy remains largely unchanged throughout training. A similar trend appears in the test loss, which remains relatively stable until a sudden bump occurs, after which the test loss settles again. This bump resembles the formation of an induction head circuit (Olsson et al., 2022), and to the best of our knowledge has previously only been observed during the training of multi-layer transformer architectures. However, as opposed to the 2-layer models, this phase transition in the loss does not correspond to an accuracy improvement for attention. Based on previous work  (Olsson et al., 2022), we hypothesize that during this phase transition, the Attention mechanism attempts to form induction heads. However a single-layer transformer lacks the expressivity needed to effectively leverage this mechanism for task resolution. Interestingly, the dynamic of Mamba is mixed:
@@
    Like single-layer Attention models, we report a significant loss bump, reinforcing the connection between Mamba and Attention mechanisms, as suggested in  Ali et al. (2024); Dao & Gu (2024).
@@
Our results highlight how Attention and recurrent models share some common ground, yet distinct inductive biases. Moreover, their performance strongly interacts with the optimization algorithm at hand (in our case, Adam (Kingma, 2014)), as we also saw in Fig. 1. Understanding these nuances is key to optimally leverage both architectures, towards hybrid models (Waleffe et al., 2024).
7 Architectural Drivers of Performance and Stability
Our results so far highlight key differences between Transformers and SSMs. Notably, while Mamba demonstrates greater expressivity – successfully solving the task even in a single-layer setting – it presents optimization challenges in terms of learning rate stability. In contrast, Transformers exhibit remarkable stability across a wide range of learning rates during training in the 2-layer setting. To address this discrepancy, we conduct a series of ablation studies aimed at:
@@ 7 Architectural Drivers of Performance and Stability
    aligning the backbone of both models (full details in Appendix A.1) and identify the source of Mamba’s superior performance in 1-layer, summarized in Table 2 and Appendix 3;
@@
Convolutions.
Inspired by  (Li et al., 2024), we begin by aligning Attention to Mamba. We incorporate a 1D convolution before the Query, Key and Value matrix projections to bring in locality, enabling the model to solve MQAR with just one layer. Interestingly, we observe that applying the convolution to either the Key or Value matrix alone is sufficient to achieve the same performance gains. These observations suggest that the 1D convolution may be a central factor behind Mamba’s effectiveness. However, we find that removing the convolution does not affect performance.
Backbone ablation.
We further modify the Mamba architecture by: (1) removing the gating mechanism, and (2) replacing the standard Mamba block with the individual sequence mixer S6, followed by an MLP – mirroring the Transformer’s architecture. Despite these alterations, Mamba performs well when properly tuned, suggesting the sequence mixer (S6) is at the root of its expressivity.
Table 2: MQAR performance of ablated 1-layer architectures. Convolutions allows Attention to reach perfect accuracy, while Mamba preserves performance in all ablations.
Model Accuracy
Attention $2\%$
Attention + Conv on QKV $99\%$
Mamba $99\%$
@@
S6 + MLP (as a Transformer) $98\%$
Positional Encodings.
We tested whether adding various Positional Embedding (PE) strategies could improve SSM performance and stability. Our findings reported in Table 4 show that PE has a negligible impact on performance. This result reinforces that the core recurrent structure is the dominant mechanism for processing sequence order in these models.
@@
To better understand what contributes to training stability, we also evaluate architectural variants designed for improving Mamba and solve the MQAR task. In particular, we test Mamba2 (Dao & Gu, 2024) and DeltaNet (Yang et al., 2024) as shown in Figure 7. While the performance of Mamba2 is slightly more stable, Transformer-level robustness is only achieved by DeltaNet. A closer look at the DeltaNet update rule reveals that its mixing is based on Householder matrices. As such, the off-diagonal terms such as $C_{N}\prod_{k=1}^{N}A_{k}B_{0}$ do not necessarily incur in vanishing gradients. Instead, in both Mamba and Mamba2, $A_{k}$ includes a decay rate that induces vanishing gradients and fast decay of off-diagonal terms, as recently pointed out by Trockman et al. (2024). We hypothesize this is the main distinction unlocking stable optimization in DeltaNet.
Figure 7: Performance of newer 1-layer models on MQAR. Here we show how having a larger hidden state marginally helps stability, as in Mamba2 and especially in DeltaNet. Our results are presented for model dimensions up to 256, which was the maximum size supported by the DeltaNet implementation. The results show the mean and relative max-min errors using 3 seeds.
@@ 8 Discussion and Conclusions
In this work, we dissected the practical differences between Transformers and modern recurrent models on the associative recall and copying tasks. Our findings demonstrate that a crucial differentiator lies not just in their theoretical expressivity, but in their fundamental learnability. We demonstrated that modern SSMs exhibit a critical optimization instability, with success confined to a narrow learning rate window—a finding that re-contextualizes prior performance evaluations. Additionally, we observed contrasting scaling behaviors: recurrent models benefit from increased width, whereas Transformers struggle in a single-layer configuration. Interestingly, despite their poor performance, single-layer Transformers exhibit training dynamics resembling the induction head phenomenon, previously reported only in multi-layer settings. Instead, Mamba displays similar behavior but successfully solves the task. Finally, our ablations show how the performance of Mamba is robust to the removal specific backbone components such as gating and convolution. More recent architectures, like DeltaNet, can enhance performance and stability.
The central implication of our work is that future research on efficient sequence models should treat optimization stability as a first-class objective, alongside expressivity and computational cost. While our findings are compelling, we acknowledge that our analysis is conducted on synthetic benchmarks highly correlated with in-context learning. Validating these dynamics on downstream language modeling tasks is a critical next step. Furthermore, a formal theoretical explanation for the optimization brittleness we empirically observe remains an important open question. Looking ahead, by showing that modern recurrent models can be as expressive as Transformers on these tasks but are harder to train, our work underscores the importance of learnability in the path towards understanding and building the next generation of sequence models.